eljaja:

60.00 grams of hydrochloric acid react with sodium hydroxide. - write the balanced reaction with the correct products - what is the reaction type? - what is the theoretical yield of sodium chloride (in grams)? using dimensional analysis - what is the percent yield of 84.05 grams of sodium chloride are actually produced?

Vocaloid:

NaOH + HCl --> switch the cations to get NaCl + HOH (better to write HOH as H2O, though) as the products

Vocaloid:

"reaction type" honestly depends on what level of chem this is, if they're talking about the movement of ions this is a double replacement (since both of them swap their cations with each other), if they're talking about what actually is happening, chemically, this is an acid-base neutralization reaction

Vocaloid:

theoretical yield of sodium chloride: start with 60.00g HCl as stated in the problem divide by the molar mass of HCl to get moles of HCl you have a 1:1 ratio of HCl and NaCl according to the equation. so moles of NaCl = moles of HCl so from there you can just multiply by molar mass of NaCl to get theoretical yield

Vocaloid:

percent yield = divide (actual yield)/(theoretical yield) and multiply by 100% don't forget sig figs on questions 3 and 4

eljaja:

thank you.. can you check my work for theoretical yield? 60.00g / 36.36 = 1.65 * 58.44 = 96.42 I might be missing a step here ^

Vocaloid:

generally you don't want to round until the very end 60.00g / 36.36 * 58.44 = 96.44g

Vocaloid:

otherwise good

eljaja:

ah okay! thanks!

eljaja:

Hi, sorry coming back to this question. For theoretical yield its right but I need to do dimensional analysis for it. Are you able to help me with it?

Vocaloid:



eljaja:

thank you!!